### v1

# Emulating Self-attention with Convolution for Efficient Image Super-Resolution

Abstract
In this paper, we tackle the high computational cost of transformers for lightweight image super-resolution (SR). Motivated by the observations of self-attention’s inter-layer repetition, we introduce a convolutionized self-attention module named Convolutional Attention (ConvAttn) that emulates self-attention’s long-range modeling capability and instance-dependent weighting with a single shared large kernel and dynamic kernels. By utilizing the ConvAttn module, we significantly reduce the reliance on self-attention and its involved memory-bound operations while maintaining the representational capability of transformers. Furthermore, we overcome the challenge of integrating flash attention into the lightweight SR regime, effectively mitigating self-attention’s inherent memory bottleneck. We scale up window size to 32$\times$32 with flash attention rather than proposing an intricated self-attention module, significantly improving PSNR by 0.31dB on Urban100$\times$2 while reducing latency and memory usage by 16$\times$ and 12.2$\times$. Building on these approaches, our proposed network, termed Emulating Self-attention with Convolution (ESC), notably improves PSNR by 0.27 dB on Urban100$\times$4 compared to HiT-SRF, reducing the latency and memory usage by 3.7$\times$ and 6.2$\times$, respectively. Extensive experiments demonstrate that our ESC maintains the ability for long-range modeling, data scalability, and the representational power of transformers despite most self-attentions being replaced by the ConvAttn module.
1 Introduction
Image Super-Resolution (SR) aims to reconstruct high-resolution (HR) images from low-resolution (LR) inputs and remains an active area of research in computer vision. Recently, the demand for multimedia content and generative models has increased significantly, making SR particularly noteworthy as it enables users to enjoy high-quality content under resource-constrained conditions. Consequently, practical deployment has emerged as a critical consideration in SR tasks, motivating numerous SR studies to enhance performance while reducing computational complexity (FLoating point OPerations; FLOPs) and parameter size.
Therefore, transformers have gained significant attention in the SR task since they have demonstrated superior performance over Convolutional Neural Networks (CNNs) while requiring lower FLOPs and fewer parameters. By capturing long-range dependencies and performing input-dependent weighting through self-attention, transformers exhibit high representational capacity that enhances performance, especially as the training data volume increases [11, 3]. Accordingly, numerous transformer-based methods have been proposed [55, 45, 37] to capitalize on the advantages of self-attention while reducing FLOPs and parameter counts.
Figure 1: Comparison of performance, latency, and memory usage on the Urban100$\times$4 dataset. Our methods are evaluated against representative SR models, including CNNs, transformers, and SSMs.
However, many of these works overlook the excessive memory access caused by self-attention, which arises from materializing the score matrix ($S=QK^{T}$[7] and leveraging memory-bound operations such as tensor reshaping and window masking [50, 31]. The memory access issue is further worsened in the SR architecture, which deals with large feature maps without patchify stem or downsampling stage. For example, SwinIR-light [28] is 4.7$\times$ slower latency and 2$\times$ higher memory usage than CNNs [57] reconstructing an HD image at the $\times$2 scale, even though it requires 14.5$\times$ and 17$\times$ fewer FLOPs and parameter sizes, respectively. Consequently, despite their promising performance, transformers are challenging to deploy on resource-constrained devices, such as consumer-level GPUs.
In this paper, we aim to design an efficiency-driven transformer tailored for lightweight SR tasks, achieving both improved performance and memory overhead. We start by conducting a preliminary analysis of self-attention, observing that the similarity modeling performed by self-attention and resulting extracted features remain highly consistent across several layers, as shown in Figure 2. This finding exhibits that self-attention may extract overlapping features, suggesting the possibility of reducing computational overhead without compromising representational capability by utilizing efficient alternatives. Building on this finding, we establish a design strategy that retains self-attention only in the first layer of each block while replacing the remaining layers with our proposed efficient alternative, the Convolutional Attention (ConvAttn) module. To effectively emulate self-attention’s long-range modeling and instance-dependent weighting, ConvAttn operates with a twofold mechanism. First, it simplifies self-attention’s long-range interaction by applying convolution with a shared 13$\times$13 large kernel throughout the entire network, targeting only a subset of channels. Second, dynamic kernels are generated to capture input-dependent weighting, mimicking the adaptive nature of self-attention. By combining these components, ConvAttn significantly reduces the reliance on memory-intensive self-attention while maintaining the representational capability of transformers.
With most self-attention layers replaced by ConvAttn, we leverage this efficiency to further enhance the remaining self-attention layers. Specifically, we enlarge the window size of self-attention, significantly improving performance with only a slight increase in FLOPs. However, increasing the window size leads to an enlarged score matrix, substantially raising peak memory usage. To address this, we introduce flash attention [7, 10] into lightweight SR tasks, to avoid materializing the score matrix. Our optimized implementation allows us to scale up the window size to 32$\times$32 while reducing latency and memory usage by 16$\times$ and 12.2$\times$, respectively, as illustrated in Table 1.
Based on these approaches, we introduce the lightweight SR network called Emulating Self-attention with Convolution (ESC). The proposed ESC outperforms ATD-light [54] on Urban100$\times$4 with 0.1dB improvements on PSNR while being 8.9$\times$ faster, as shown in Figure 1. Additionally, ESC-light surpasses ELAN-light [55] by a large margin of 0.29dB in PSNR on Urban100$\times$2 while also reducing latency by 22%. We further validate our ESC in scenarios where reducing FLOPs and parameter size is essential by introducing ESC-FP, which outperforms MambaIRV2-light [17] on Manga109$\times$4 with reductions of FLOPs and parameter size by 20% and 32%, respectively. Through our extensive experiments, we demonstrate that our ESC fully leverages the advantages of transformers – including their large receptive fields, representational capacity, and scalability with respect to data volume – even though most self-attentions are replaced by the ConvAttn module. We support these results with in-depth experiments, as illustrated in Figure 6, suggesting that the proposed ConvAttn module extracts similar features with self-attention.
Our contributions are summarized as follows:
  • We demonstrate that well-designed convolution partially replaces self-attention, significantly improving efficiency without sacrificing the advantages of transformers.
  • We mitigate the memory overhead of self-attention by integrating flash attention, thereby enlarging window size without excessive memory usage. To the best of our knowledge, this is the first successful application of flash attention in lightweight SR tasks.
  • We fully exploit the transformer’s benefits in lightweight SR tasks by making them more efficient and simpler.
2 Related Work
Traditional CNNs Early deep learning‐based SR research [9, 22, 29, 58, 57] primarily employed convolution that extracts local features from images. These approaches typically relied on stacked convolutions, which often resulted in over‐parameterization. To address this issue, numerous studies [21, 42, 27, 36] have shared their weights across multiple layers to reduce parameter size. However, many of them suffer from performance degradation since the input-independent weights are shared across multiple layers. In this paper, we share a large kernel through the entire network while leveraging input-dependently generated dynamic DWC [18] by each layer in parallel. This approach efficiently mitigates the parameter growth of the large kernel while maintaining each layer’s representational power.
Transformers Pioneering studies that significantly reduce computational costs by computing self-attention within local windows [28] or channel-wise [51] have sparked considerable interest in applying transformers [44, 11] to the SR task. Subsequent studies have proposed methods that enhance the self‐attention mechanism of transformers to strengthen their long‐range modeling capabilities while reducing both FLOPs and parameters [55, 45, 5, 62, 54, 56, 37]. Nonetheless, transformers remain challenging to deploy due to the excessive memory access caused by self-attention and involved memory-bound operations. In this paper, we address this limitation by replacing self‐attention with convolutions that emulate its advantages, thereby preserving the advantages of transformers while reducing reliance on self-attentions and associated memory-bound operations.
Large Kernel CNNs and State-Space Models Several recent studies have highlighted the computational burden of transformers and investigated alternative core operators to replicate their advantages. Specifically, CNN-based models [41, 61, 49] attempt to mimic the advantages of transformers by leveraging large kernels and pixel attention [59], while State-Space Models (SSM) [16, 17] adapt their verified long-range modeling abilities from 1D sequence domain [14, 13] to the 2D image domain. However, these approaches often overlook either the exceptional representational capability brought by self-attention’s sophisticated similarity modeling or the additional memory access caused by supplementary mechanisms to adapting 2D domains like multi-directional or attentive scanning. In this paper, we fully leverage the advantages of both self-attention and convolution while utilizing flash attention [7, 10] to address the memory access of self-attention without relying on complex additional mechanisms.
Figure 2: Visualized inter-layer similarities of the feature extracted by self-attention ($F$) before added to skip path and corresponding attention map ($A^{avg}$). Both CKA [24] and cosine similarities are measured on the Urban100$\times$4 dataset. The window and head dimensions of the attention map are averaged for visualization.
3 Proposed Methods
In this section, we start by representing our preliminary analysis and then describe our proposed network.
3.1 Preliminary analysis
Transformers often achieve improved performance by utilizing self-attention’s long-range modeling capabilities and instance-dependent weighting. However, each self-attention incurs excessive memory overhead, resulting in high latency and memory usage. Interestingly, our preliminary research demonstrates that the sophisticated similarity modeling of self-attention ($A^{avg}$) and the resulting extracted features ($F$) exhibit a high degree of inter-layer similarities, with averages of 89% and 87%. This observation not only aligns with recent research [60] but also explains the success of studies that reduced computational cost by sharing the attention map across layers [55, 37]. Inspired by this observation, we investigate improving efficiency without degrading the representational capability by replacing self-attention with efficient alternatives.
3.2 Overall Structure
Figure 3: Visualization of the overall architecture.
Our network is composed of four main components, as shown in Figure 3. First, the LR input image $I_{LR}\in\mathbb{R}^{H\times W\times 3}$ is transformed into a shallow feature $F_{0}\in\mathbb{R}^{H\times W\times C}$ through a 3$\times$3 convolution ($\mathrm{Conv}_{3\times 3}$). Here, $H$ and $W$ denote height and width $I_{LR}$, and $C$ denotes feature dimension. Next, $F_{0}$ and a large kernel $LK\in\mathbb{R}^{13\times 13\times 16\times 16}$ that is shared throughout the entire model are fed into the deep feature extractor $H$, and the resulting deep feature is added back to $F_{0}$ to produce the final feature $F\in\mathbb{R}^{H\times W\times C}$. In parallel, the image-wise skip module $S$ takes $I_{LR}$ as input to generate a skipped feature $F_{s}$, and the upsampler $U$ utilize both $F$ and $F_{s}$ as inputs to return the upsampled image $I_{SR}\in\mathbb{R}^{rH\times rW\times 3}$ where $r$ denotes upscaling factor. The entire process is formulated as follows:
$F_{0}$ $=\mathrm{Conv}_{3\times 3}(I_{LR}),$ (1)
$F$ $=H(F_{0},LK)+F_{0},$
$F_{s}$ $=S(I_{LR}),$
$I_{SR}$ $=U(F,F_{s}).$
The $H$ is composed of the $N$ $\mathrm{ESCBlock}$s where $N$ denotes the number of $\mathrm{ESCBlock}$s and a 3$\times$3 convolution, and is formulated as follows:
$F_{i}$ $=\mathrm{ESCBlock}_{i}(F_{i-1},LK),\quad i=1,\dots,N,$ (2)
$F_{d}$ $=\mathrm{Conv}_{3\times 3}(F_{N}),$
where $F_{i-1}$ indicates the feature input to the $i$th $\mathrm{ESCBlock}$ together with $LK$, and $F_{i}$ represents the extracted feature. The $U$ and $S$ will be described in the implementation detail section because they will be changed depending on the task.
3.3 ESC Block
The overall process of the $\mathrm{ESCBlock}$ is illustrated in Figure 4 and is formulated as follows:
$F^{in}_{i}$ $=\mathrm{ConvFFN}\bigl{(}\mathrm{LN}(F_{i-1})\bigr{)},$ (3)
$F_{i,0}$ $=F^{in}_{i}+\mathrm{SelfAttn}\bigl{(}\mathrm{LN}(F^{in}_{i})\bigr{)},$
$F_{i,j}$ $=F_{i,j-1}+\mathrm{ConvAttn}_{j}\bigl{(}\mathrm{ConvFFN}_{j}(F_{i,j-1}),LK\bigr{)},$
$\quad\quad j=1,\dots,M,$
$F_{i}$ $=F_{i-1}+\mathrm{Conv}_{3\times 3}\bigl{(}\mathrm{LN}(F_{i,M})\bigr{)}.$
Firstly, the $F_{i-1}$ is converted to the $F^{in}_{i}$ by layer normalization ($\mathrm{LN}$) and convolutional FFN ($\mathrm{ConvFFN}$[62] with kernel size three. Then, $\mathrm{LN}$ and self-attention ($\mathrm{SelfAttn}$) with a window size of 32 is employed to residual connection to extract $F_{i,0}$. Despite the large window size, we effectively reduce memory overhead by utilizing flash attention. Furthermore, since ConvFFN is placed before, the self-attention extract features considering local information without the complicated $QKV$ projection module. After establishing $F_{i,0}$, the block iterates residual paths composed of $\mathrm{ConvFFN}$ and proposed ConvAttn module ($\mathrm{ConvAttn}$) $M$ times. Although our methods do not utilize shifted-window self-attention, these convolution-based modules, after the self-attention, will extract inter-window features. Finally, the result of the last iteration $F_{i,M}$ go through $\mathrm{LN}$ and $\mathrm{Conv}_{3\times 3}$ and is added to $F_{i-1}$, resulting in $F_{i}$.
Figure 4: Visualization of the proposed ESC block.
Convolutional Attention Module The $\mathrm{ConvAttn}$ is designed to emulate self-attention’s two key advantages: capturing long-range dependencies and instance-dependent weighting. For this purpose, $\mathrm{ConvAttn}$ leverages to types of convolution formulated as:
$F^{att}_{i,j},F^{idt}_{i,j}$ $=\mathrm{Split}_{16:C-16}(F^{\mathrm{CF}}_{i,j}),$ (4)
$DK_{i,j}$ $=\mathrm{Conv}^{up}_{1\times 1}\bigl{(}\phi(\mathrm{Conv}^{down}_{1\times 1}(\mathrm{GAP}(F^{att}_{i,j})))\bigr{)},$
$F^{res}_{i,j}$ $=(F^{att}_{i,j}\circledast DK_{j})+(F^{att}_{i,j}\circledast LK),$
$F^{fuse}_{i,j}$ $=\mathrm{Conv}^{fuse}_{1\times 1}\bigl{(}\mathrm{Concat}(F^{fuse}_{i,j},F^{idt}_{i,j})\bigr{)},$
where $F^{\mathrm{CF}}_{i,j}$ denotes the features extracted by $\mathrm{ConvFFN}_{i,j}$ before. We first split the $F^{\mathrm{CF}}_{i,j}$ channel-wise, resulting in $F^{att}_{i,j}\in\mathbb{R}^{H\times W\times 16}$, which consists of the 16 channels in front, and $F^{idt}_{i,j}\in\mathbb{R}^{H\times W\times C-16}$, which contains the remaining part. Both large kernel and dynamic convolutions are only performed within $F^{att}_{i,j}$ to reduce memory access cost, unlike the methods that perform large kernel convolution to full channels [41, 49]. Next, dynamic depth-wise kernel $DK_{i,j}\in\mathbb{R}^{3\times 3\times 1\times 16}$ [18] is created by two layers of $\mathrm{Conv}_{1\times 1}$ with global average pooling ($\mathrm{GAP}$) and gaussian error linear unit activation ($\phi$). Since enlarging $DK$’s kernel size increases parameter size multiplied by the hidden dimension of the kernel estimator, we fix its kernel size to 3$\times$3. Then both $DK_{i,j}$ and $LK$ are convolved on $F^{att}_{i,j}$ and sum up those, resulting convolved feature $F^{res}_{i,j}$. Inspired by previous studies [55, 37] that shared attention maps across multiple layers and our preliminary analysis, we hypothesize that long-range interactions may not significantly change with varying layers. Therefore, we share $LK$ throughout the entire model, mitigating parameter growth and corresponding optimization difficulty [30, 4]. Consequently, harmonizing both $LK$ and $DK_{i,j}$ allows us to emulate self-attention’s advantages of capturing long-range dependencies and performing instance-dependent weighting with minimal increase in latency and parameter size. Finally, $F^{res}_{i,j}$ and $F^{idt}_{i,j}$ are concatenated channel-wise, followed by the last $\mathrm{Conv}_{1\times 1}$ to fuse them, resulting in fused feature $F^{fuse}_{i,j}$. In this way, both the local features extracted by $\mathrm{ConvFFN}$ and the long-range structural features extracted by $LK$ and $DK$ can be leveraged to capture multi-scale features.
While $DK_{i,j}$ can be structurally re-parameterized into $LK$ by padding both spatial and channel-wise to reduce FLOPs, we do not because it slows down latency. On the other hand, in situations where reducing FLOPs and parameter size is crucial rather than reducing latency, $LK$ can be decomposed depth-wise separable manner [19] into point-wise large kernel filter $LK^{\mathrm{c}}\in\mathbb{R}^{1\times 1\times 16\times 16}$ and depth-wise large kernel filter $LK^{\mathrm{s}}\in\mathbb{R}^{13\times 13\times 1\times 16}$. This decomposition redefines line 3 in Equation 4 as follows:
$F^{res}_{i,j}=(F^{att}_{i,j}\circledast LK^{c})\circledast\bigr{(}ZP(DK_{i,j})+LK^{s}\bigl{)},$ (5)
where $ZP(\cdot):\mathbb{R}^{3\times 3\times 1\times 16}\mapsto\mathbb{R}^{13\times 13\times 1\times 16}$ denotes spatial zero padding.
In summary, the ConvAttn module efficiently emulates self-attention’s two advantages by leveraging a shared $LK$ and layer-specific $DK$s. By replacing self-attention and associated memory-bound operations with the proposed ConvAttn module, we significantly improve efficiency while maintaining the benefits of transformers.
Table 1: Comparison of latency and memory usage between naive self-attention and flash attention (FA) [7, 10]. WS and #Pixels denote the window size of self-attention and the corresponding number of pixels. $X$ refers to the input tensor.
WS (#Pixels) Type $X\in\mathbb{R}^{64\times 64\times 64}$ $X\in\mathbb{R}^{288\times 288\times 64}$
Latency
Memory Usage
Latency
Memory Usage
16$\times$16 (256) w/o FA 2.0 46 4.9 758
w/ FA 0.5 (4×4$\times$4 ×) 11 (4.2×4.2$\times$4.2 ×) 1.3 (3.8×3.8$\times$3.8 ×) 223 (3.4×3.4$\times$3.4 ×)
24$\times$24 (576) w/o FA 7.6 111 14.5 1575
w/ FA 0.5 (15.2×15.2$\times$15.2 ×) 14 (7.9×7.9$\times$7.9 ×) 2.0 (7.3×7.3$\times$7.3 ×) 223 (7.1×7.1$\times$7.1 ×)
32$\times$32 (1024) w/o FA 27.3 157 44.7 2717
w/ FA 0.5 (54.6×54.6$\times$54.6 ×) 11 (14.3×14.3$\times$14.3 ×) 2.8 (16.0×16.0$\times$16.0 ×) 223 (12.2×12.2$\times$12.2 ×)
Table 2: Comparisons of classic SR methods trained on the DIV2K dataset. Mem, #FLOPs, and #Params denote memory usage, the number of FLOPs, and parameter size, respectively. $\dagger$ denotes that we re-calculate the statistics of the indicated methods. All statistics and metrics are measured as detailed in Section 4.2. The best result on each PSNR and SSIM is bolded.
Method Scale Latency (ms) Mem (mb) #FLOPs (G) #Params (K) PSNR / SSIM
Set5 Set14 B100 Urban100 Manga109
SwinIR-lt [28] ×$\times 2$ 1409.8 1287 244.2 910 38.14/0.9611 33.86/0.9206 32.31/0.9012 32.76/0.9340 39.12/0.9783
ELAN-lt [55] 94.5 887 203.1 621 38.17/0.9611 33.94/0.9207 32.30/0.9012 32.76/0.9340 39.11/0.9782
OmniSR [45] 120.3 1031 194.5 772 38.22/0.9613 33.98/0.9210 32.36/0.9020 33.05/0.9363 39.28/0.9784
SRFormer-lt [62] 1456.3 1184 236.3 853 38.23/0.9613 33.94/0.9209 32.36/0.9019 32.91/0.9353 39.28/0.9785
ATD-lt [54] 733.5 2839 380.0 753 38.29/0.9616 34.10/0.9217 32.39/0.9023 33.27/0.9375 39.52/0.9789
HiT-SRF [56] 268.1 1804 226.5 847 38.26/0.9615 34.01/0.9214 32.37/0.9023 33.13/0.9372 39.47/0.9787
ASID-D8 [37] 131.2 999 190.5 732 38.32/0.9618 34.24/0.9232 32.40/0.9028 33.35/0.9387 -/-
MambaIR-lt [16] 277.1 1695 334.2 905 38.13/0.9610 33.95/0.9208 32.31/0.9013 32.85/0.9349 39.20/0.9782
MambaIRV2-lt [17] 580.4 2824 286.3 774 38.26/0.9615 34.09/0.9221 32.36/0.9019 33.26/0.9378 39.35/0.9785
RDN [58] 279.3 2058 5096.2 22123 38.24/0.9614 34.01/0.9212 32.34/0.9017 32.89/0.9353 39.18/0.9780
RCAN [57] 299.3 626 3529.7 15445 38.27/0.9614 34.12/0.9216 32.41/0.9027 33.34/0.9384 39.44/0.9786
ESC-FP (Ours) 94.3 627 239.8 524 38.27/0.9617 34.09/0.9219 32.37/0.9022 33.22/0.9375 39.40/0.9784
ESC-lt (Ours) 73.2 830 359.4 603 38.24/0.9615 33.98/0.9211 32.35/0.9020 33.05/0.9363 39.33/0.9786
ESC (Ours) 120.9 831 592.0 947 38.35/0.9619 34.11/0.9223 32.41/0.9027 33.46/0.9395 39.54/0.9790
SwinIR-lt [28] ×$\times 3$ 331.7 596 110.8 918 34.62/0.9289 30.54/0.8463 29.20/0.8082 28.66/0.8624 33.98/0.9478
ELAN-lt [55] 32.5 399 90.1 629 34.61/0.9288 30.55/0.8463 29.21/0.8081 28.69/0.8624 34.00/0.9478
OmniSR [45] 41.2 476 88.4 780 34.70/0.9294 30.57/0.8469 29.28/0.8094 28.84/0.8656 34.22/0.9487
SRFormer-lt [62] 530.5 537 105.4 861 34.67/0.9296 30.57/0.8469 29.26/0.8099 28.81/0.8655 34.19/0.9489
ATD-lt [54] 274.4 1258 168.0 760 34.74/0.9300 30.68/0.8485 29.32/0.8109 29.17/0.8709 34.60/0.9506
HiT-SRF [56] 124.9 1464 101.6 855 34.75/0.9300 30.61/0.8475 29.29/0.8106 28.99/0.8687 34.53/0.9502
ASID-D8 [37] 61.9 460 86.4 739 34.84/0.9307 30.66/0.8491 29.32/0.8119 29.08/0.8706 -/-
MambaIR-lt [16] 109.3 760 148.5 913 34.63/0.9288 30.54/0.8459 29.23/0.8084 28.70/0.8631 34.12/0.9479
MambaIRV2-lt [17] 259.0 1250 126.7 781 34.71/0.9298 30.68/0.8483 29.26/0.8098 29.01/0.8689 34.41/0.9497
RDN [58] 146.5 985 2281.2 22308 34.71/0.9296 30.57/0.8468 29.26/0.8093 28.80/0.8653 34.13/0.9484
RCAN [57] 85.1 560 1586.1 15629 34.74/0.9299 30.65/0.8482 29.32/0.8111 29.09/0.8702 34.44/0.9499
ESC-FP (Ours) 34.8 291 110.0 530 34.72/0.9300 30.67/0.8483 29.30/0.8107 29.12/0.8706 34.56/0.8706
ESC-lt (Ours) 25.2 384 162.8 612 34.61/0.9295 30.52/0.8475 29.26/0.8102 28.93/0.8679 34.33/0.9495
ESC (Ours) 41.4 385 267.6 955 34.84/0.9308 30.74/0.8493 29.34/0.8118 29.28/0.8739 34.66/0.9512
SwinIR-lt [28] ×$\times 4$ 222.9 351 63.6 930 32.44/0.8976 28.77/0.7858 27.69/0.7406 26.47/0.7980 30.92/0.9151
ELAN-lt [55] 18.0 241 54.1 640 32.43/0.8975 28.78/0.7858 27.69/0.7406 26.54/0.7982 30.92/0.9150
OmniSR [45] 22.5 273 50.9 792 32.49/0.8988 28.78/0.7859 27.71/0.7415 26.64/0.8018 31.02/0.9151
SRFormer-lt [62] 287.2 329 62.8 873 32.51/0.8988 28.82/0.7872 27.73/0.7422 26.67/0.8032 31.17/0.9165
ATD-lt [54] 189.7 753 100.1 769 32.63/0.8998 28.89/0.7886 27.79/0.7440 26.97/0.8107 31.48/0.9198
HiT-SRF [56] 82.1 1331 58.0 866 32.55/0.8999 28.87/0.7880 27.75/0.7432 26.80/0.8069 31.26/0.9171
ASID-D8 [37] 61.8 265 49.6 748 32.57/0.8990 28.89/0.7898 27.78/0.7449 26.89/0.8096 -/-
MambaIR-lt [16] 55.8 438 84.6 924 32.42/0.8977 28.74/0.7847 27.68/0.7400 26.52/0.7983 30.94/0.9135
MambaIRV2-lt [17] 153.4 748 75.6 790 32.51/0.8992 28.84/0.7878 27.75/0.7426 26.82/0.8079 31.24/0.9182
RDN [58] 66.0 791 1309.2 22271 32.47/0.8990 28.81/0.7871 27.72/0.7419 26.61/0.8028 31.00/0.9151
RCAN [58] 52.2 540 917.6 15592 32.63/0.9002 28.87/0.7889 27.77/0.7436 26.82/0.8087 31.22/0.9173
ESC-FP (Ours) 21.7 158 60.8 539 32.56/0.8897 28.87/0.7889 27.75/0.7435 26.90/0.8098 31.40/0.9192
ESC-lt (Ours) 13.3 213 91.0 624 32.52/0.8995 28.87/0.7878 27.72/0.7423 26.76/0.8058 31.26/0.9173
ESC (Ours) 21.9 215 149.2 968 32.68/0.9011 28.93/0.7902 27.80/0.7447 27.07/0.8144 31.54/0.9207
4 Experiments
This section describes implementation details for flash attention and our extensive experiments on various tasks. For implementation details for proposed networks, please refer to Section 6 in the supplementary.
4.1 Implementation Details for Flash Attention
We aim to address the excessive memory overhead of self-attention by leveraging flash attention [7]. However, the existing flash attention implementation does not support relative positional bias [39], which is essential for performance improvements. Accordingly, we implement Flash attention leveraging flex attention [10] to incorporate relative positional bias and accelerate self-attention during inference. As shown in Table 1, our implementation significantly reduces latency and memory usage of self-attention by 16$\times$ and 12.2$\times$, respectively. As the patch size for training is 64$\times$64, we do not consider the window size above 32 as it may cause training instability due to the necessity of padding.
4.2 Classic SR Results
To evaluate our proposed networks on classic SR task, we leverage five commonly used datasets (Set5 [2], Set14 [52], B100 [33], Urban100 [20], Manga109 [34]). We measure peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) on the Y channel after cropping boundary corresponding to the upscaling factor and converting them into YCbCr space. All statistics are measured by reconstructing an HD image (1280$\times$720) image using RTX4090 GPU at FP32 precision 100 times. Memory usage is calculated by torch.cuda.max_memory_allocated function and FLOPs are measured by fvcore library. For the classic SR task, we present three variants: ESC, ESC-light, and ESC-FP. ESC and ESC-light prioritize to reduce latency, while ESC-FP focuses on reducing FLOPs and parameter size.
4.2.1 Quantitative Results on the DIV2K dataset
We compare our methods with various representative lightweight SR methods, including CNN-based models [58, 57], transformers [28, 55, 45, 62, 54, 56, 37], and state-space models [16, 17] trained on DIV2K dataset [1]. Notably, as shown in Table 2, our ESC achieves notable 0.1dB PSNR improvement on the Urban100$\times$4 dataset, while being 8.7$\times$ faster than ATD-light. Moreover, ESC-light is 1.4$\times$ faster than ELAN-light, exhibiting an impressive performance improvement of 0.34dB PSNR in the Manga109×$\times 4$ dataset. These results exhibit that state-of-the-art performance can be achieved even when most self-attentions are replaced with convolutions, suggesting replacing self-attention with convolutions does not compromise representational power. Furthermore, ESC-FP outperforms MambaIRV2-light over 0.18dB on the Urban100×$\times 4$ dataset while reducing FLOPs and parameters by 20% and 31%, respectively. This result underscores that our design is effective even when reducing FLOPs and parameter size.
4.2.2 Visual Results on the DIV2K dataset
We also visually compare our ESC with representative lightweight SR methods, including SRFormer-light [62], HiT-SRF [56], ATD-light [54], and MambaIRV2-light [17]. As shown in Figure 5, Our ESC effectively restores details and achieves a high PSNR, demonstrating that our network also produces visually pleasant results.
Figure 5: Visual comparisons for classic SR ($\times$2). The best result on PSNR is bolded.
4.2.3 Ablation Study
We conduct an ablation study to justify the proposed network’s composition, as shown in Table 3. First, we introduce two variants, one that uses only the self-attention module and another that uses only the ConvAttn module. Both variants exhibit a performance drop compared to ESC, suggesting that leveraging both self-attention and ConvAttn to balance performance and efficiency is crucial. Next, we vary the $LK$’s kernel size to 9$\times$9 and 17$\times$17 to determine the optimal size. In both cases, we observe a drop in performance, confirming that a kernel size of 13$\times$13 is optimal. Then, we evaluate performance changes by altering the convolution type used in the ConvAttn module. We observe a drop in performance in variants that neither share $LK$ nor employ input-dependent $DK$. This indicates that emulating the input-dependent weighting of self-attention is essential. Accordingly, we add $DK$ without sharing $LK$ and observe some performance improvement, though it still falls short of ESC. These results suggest that employing highly sophisticated modeling may not be necessary when using convolution to emulate self-attention’s ability to capture long-range dependencies. Instead, sharing $LK$, which reduces the number of parameters and thereby eases optimization [8, 30], appears to have a more significant impact on performance. Finally, we investigate the effect of the self-attention window size by setting it to 16 and 24 and adding extra layers. In every case, we observe a performance drop, which confirms that enlarging the window size to enhance self-attention is effective even when only a single self-attention is leveraged in each block.
Table 3: Ablation studies on our ESC network. All comparisons are designed to have similar latencies to ESC. LK, DK, and WS denote large kernel, dynamic kernel, and window size, respectively. All statistics and metrics are measured as detailed in Section 4.2.
Cases
Latency
Memory
#FLOPs
#Params
Set5$\times$2 Urban100$\times$2
Only Self-attentions 128.2 830 456.6 737 38.27/0.9617 33.23/0.9378
Only ConvAttns 126.3 514 962.4 1210 38.18/0.9613 32.91/0.9353
9$\times$9 LK 117.1 831 462.2 924 38.33/0.9618 33.42/0.9392
17$\times$17 LK 126.6 832 768.9 977 38.32/0.9618 33.40/0.9390
w/o LK Share and DK 119.7 835 591.1 1950 38.32/0.9618 33.37/0.9389
w/o LK Share 120.9 835 592.0 1985 38.32/0.9619 33.43/0.9392
WS16 and more layers 118.0 807 603.6 1054 38.24/0.9614 33.05/0.9360
WS24 and more layers 116.5 801 583.5 995 38.33/0.9618 33.28/0.9383
ESC (Ours) 120.9 831 592.0 947 38.35/0.9619 33.46/0.9395
Table 4: Comparisons of classic SR methods trained on the large-scale DFLIP dataset. Mem, #FLOPs, and #Params denote memory usage, the number of FLOPs, and parameter size, respectively. §$\S$ denotes that we train the indicated methods following their descriptions. All statistics and metrics are measured as detailed in Section 4.2. The best result on each PSNR and SSIM is bolded.
Method Scale Latency (ms) Mem (mb) #FLOPs (G) #Params (K) PSNR / SSIM
Set5 Set14 B100 Urban100 Manga109
SRFormer-lt§ [62] ×$\times 2$ 1838.1 1184 236.3 853 38.24/0.9615 34.13/0.9218 32.42/0.9026 33.37/0.9386 39.36/0.9787
ATD-lt§ [54] 733.5 2839 380.0 753 38.29/0.9616 34.30/0.9230 32.43/0.9027 33.62/0.9401 39.60/0.9791
HiT-SRF§ [56] 268.1 1804 226.5 847 38.31/0.9616 34.31/0.9230 32.45/0.9031 33.58/0.9404 39.69/0.9793
ESC (Ours) 120.9 831 592.0 947 38.34/0.9618 34.42/0.9235 32.50/0.9036 33.86/0.9424 39.73/0.9795
SRFormer-lt§ [62] ×$\times 3$ 668.3 537 105.4 861 34.67/0.9297 30.75/0.8484 29.30/0.8108 29.10/0.8701 34.26/0.9498
ATD-lt§ [54] 274.4 1258 168.0 760 34.71/0.9300 30.77/0.8493 29.33/0.8116 29.42/0.8743 34.61/0.9509
HiT-SRF§ [56] 124.9 1464 101.6 855 34.69/0.9298 30.81/0.8493 29.32/0.8115 29.28/0.8729 34.72/0.9511
ESC (Ours) 41.4 385 267.6 955 34.85/0.9312 30.97/0.8511 29.41/0.8135 29.70/0.8799 34.94/0.9525
SRFormer-lt§ [62] ×$\times 4$ 327.8 329 62.8 873 32.49/0.8993 28.89/0.7887 27.76/0.7429 26.90/0.8086 31.25/0.9189
ATD-lt§ [54] 189.7 753 100.1 769 32.52/0.8995 28.93/0.7896 27.79/0.7443 27.18/0.8150 31.47/0.9208
HiT-SRF§ [56] 82.1 1331 58.0 866 32.55/0.8997 28.96/0.7897 27.77/0.7443 27.07/0.8130 31.59/0.9208
ESC (Ours) 21.9 215 149.2 968 32.79/0.9025 29.06/0.7927 27.85/0.7466 27.45/0.8229 31.87/0.9239
4.2.4 Analysis on ConvAttn Module
To verify the proposed ConvAttn module extracts similar features as self-attention, we examine the features extracted by ConvAttn modules. Self-attentions are known to extract shape-biased structural features [38, 8], unlike 3$\times$3 convolution, which mainly extracts local texture and edges. Indeed, as shown in Figure 6, the visualized average feature maps indicate that $F^{idt}$, which passed through only the ConvFFN module with a kernel size of 3$\times$3, primarily captures textures such as hair. In contrast, the feature maps extracted using both $LK$ and $DK$ exhibit a bias towards depth and structure, indicating that the proposed ConvAttn module extracts features similar to self-attention. Additionally, we observe that the inter-layer similarity is reduced when both $LK$ and $DK$ are used, compared to using $LK$ alone This observation indicates that $DK$s are extracting diverse layer-specific features, providing robust evidence of the effectiveness of dynamic weights.
Figure 6: Visualized average feature map and inter-layer CKA similarity on the features extracted by ConvAttn. We average all ConvAttn’s outputs while restoring the Lenna in the Set14 dataset. CKA similarities are measured on the Urban100$\times$4 dataset.
4.2.5 Local Attribution Map
We use the Local Attribution Map (LAM) [15] to verify our networks can extract long-range dependencies despite using less self-attention. We compare our network with studies that aim to expand the receptive field of self-attention by using various self-attention implementations, including global-level attention [45], permuted self-attention [62], and spatial-channel correlation (which includes both window and channel-wise self-attention) [56]. Notably, as described in Figure 7, our ESC references the widest range to restore the red bounding box along with the highest diffusion index. This result exhibits our network effectively captures long-range dependencies even though most self-attentions are replaced by the ConvAttn module, suggesting that our ConvAttn module contributes to retaining the large receptive field of transformers.
Figure 7: Comparisons of Local Attribution Map (LAM) [15] and the corresponding area of contribution and diffusion index. The best result on the diffusion index is bolded.
4.2.6 Quantitaive Results on the DFLIP dataset
Transformers exhibit significant performance improvements when trained on large-scale datasets [11, 3]. To verify our model also retains this benefit, we train and evaluate various representative transformers – including SRFormer-light [62], ATD-light [54], and HiT-SRF [56] – alongside our ESC on a large dataset. For this purpose, we leverage DIV2K+Flicker2K+LSDIR+DiverSeg-IP (DFLIP) [1, 43, 26, 35] dataset. As illustrated in Table 4, our approach outperforms the other methods by a substantial margin across all scales and evaluation datasets. These results indicate that replacing self-attention with ConvAttn does not sacrifice the data scaling capability of transformers.
4.3 Arbitrary-Scale SR
To evaluate our proposed networks on arbitrary-scale SR tasks, we leverage the DIV2K-validation [1] dataset. Following previous research [6, 25], we assess PSNR in the RGB channel after cropping the boundaries according to the upscaling factor plus six. For arbitrary-scale SR tasks, we change $U$ of our network and representative lightweight SR transforemrs [54, 56] into LTE [25].
4.3.1 Quantitative Results
To demonstrate the outstanding representational capability of the proposed network, we train and evaluate the network on an arbitrary-scale SR task, which requires processing various upscaling factors with a single network. For quantitative comparison, we leverage commonly used RDN [58] and additional representative transformers [54, 56] that we train. As described in Tables 5, our network continually achieves high performance at the scale both seen and unseen during training, even achieving a 0.09dB improvement over RDN+LTE at $\times$2. This result underscores our ESC’s exceptional representational capability. For quantitative results measured on other datasets, including Set5, Set14, B100, and Urban100, refer to Section 7 in the supplementary.
Table 5: Quantitative comparison on arbitrary-scale SR task employing LTE [25] as upsampler. The best result on PSNR is bolded.
Methods Seen Unseen
$\times$2 $\times$3 $\times$4 $\times$6 $\times$12 $\times$18 $\times$24 $\times$30
RDN+LTE [58] 35.04 31.32 29.33 27.04 23.95 22.40 21.36 20.64
ATD-lt+LTE§ [54] 34.98 31.25 29.27 27.01 23.94 22.38 21.37 20.64
HiT-SRF+LTE§ [56] 35.10 31.35 29.35 27.05 23.94 22.39 21.36 20.64
ESC+LTE (Ours) 35.13 31.38 29.37 27.08 23.97 22.40 21.38 20.64
4.3.2 Visual Results
We also visually compare our network with RDN, HiT-SRF, and ATD-light. As shown in Figure 8, Our network effectively restores fine details that other methods fail on both $\times$4 and $\times$12 scales. This result demonstrates that our network also excels at restoring visually present results on both in-distribution and out-of-distribution scales.
Figure 8: Visual comparisons on arbitrary-scale SR for both in-distribution ($\times$4) and out-of-distribution ($\times$12) scales. The best result on PSNR is bolded.
4.4 Real-World Super-Resolution
We further test our network on real-world SR tasks that aim to restore images with unknown degradations. To this end, we introduce a variant called ESC-Real and compare them to representative real-world SR methods, including RealESRGAN [47], SwinIR-Real [28], and DASR [48]. For our evaluation, we use the commonly employed RealSRSet dataset [53]. As shown in Table 9, our ESC-Real recovers more fine-grained details than the other methods while significantly reducing unpleasant artifacts in both graphic and natural images. This result demonstrates that our proposed methods are effective even when dealing with complex unknown degradations, underscoring our methods’ outstanding representational capability.
Figure 9: Visual comparisons on real-world SR ($\times$4) tasks.
5 Conclusion
In this paper, we propose ESC, a novel lightweight SR network designed to mitigate the excessive memory overhead of transformers. Our extensive experiments demonstrate that most self-attention layers can be effectively replaced with carefully designed convolution to improve efficiency. Moreover, we introduce flash attention into lightweight SR tasks, significantly enhancing performance without complicated modifications to self-attention. As a result, we fully exploit the advantages of transformers by prioritizing improvements in their efficiency.
Supplementary Material
The supplementary includes implementation details of our proposed networks and additional results on the arbitrary-scale super-resolution (SR) tasks.
6 Implementation Details
This section describes the details of the implementation of our methods. We begin by outlining the basic model configuration for each task, which includes the hyper-parameters $C$, $N$, and $M$. Here, $C$ represents the number of channels, $N$ denotes the number of $\mathrm{ESCBlock}$s, and $M$ indicates the number of $\mathrm{ConvAttn}$s. Next, we discuss additional components, such as the compositions of $U$ and $S$, among others. Finally, we present the training details, which cover the training datasets, the number of training iterations, the learning rate, optimizer configurations, the loss function used, and the training batch and patch size.
6.1 Classic SR
Three variants are introduced in the classic SR task: ESC-FP, ESC-light, and ESC. ESC-FP is a variant needed when it is necessary to reduce FLOPs and parameter size, while ESC-light and ESC are variants needed when it is necessary to reduce latency. The basic configurations for each variant are as follows: for ESC-FP, $C$, $N$, and $M$ are 48, 5, and 5; for ESC-light, they are 64, 3, and 5; and for ESC, they are 64, 5, and 5. All variants use Sub-Pixel Convolution (SPConv) [40] as $U$. ESC-light and ESC utilize the repeat function [12] as $S$ and add $F^{s}$ before the pixel shuffle of SPConv, while ESC-FP employs bicubic interpolation as $S$ and adds $F^{s}$ to the output of SPConv. Furthermore, ESC-FP employs decomposed $LK$ to further reduce FLOPs and parameter size and utilizes extra layer normalization in front of the $\mathrm{ConvFFN}$s. For training, we use the DIV2K [1] dataset, and for the data scaling experiments, we employ the DIV2K+Flicker2K+LSDIR+Diverseg-IP (DFLIP) [43, 26, 35] dataset. Training our networks lasts for 500K iterations, and the optimizer used is AdamW [32] with $\beta$s of 0.9 and 0.9 and a learning rate of 5e-4. We use L1 loss and 64 patches of size 64$\times$64 as input to train. The networks of scale ×$\times 3$ and $\times$4 are fine-tuned from the results of the $\times$2 scale. To train other methods [62, 56, 54] for data scaling experiments, we follow the training details described in their paper.
6.2 Arbitrary-scale SR
For the Arbitrary-scale SR task, we use the same basic configuration as the ESC. The difference between the ESC in Classic SR and the ESC in arbitrary-scale SR is that LTE [25] is used as $U$, and accordingly, $S$ is also changed to bilinear interpolation. Training details, including other models [54, 56], are the same as LTE across all instances, using the DIV2K dataset, running for 1000 epochs, utilizing the Adam [23] with $\beta$s values of 0.9 and 0.999, and leveraging L1 loss. However, since HiT-SRF and our ESC are optimized for training with the input patches of 64$\times$64, we train all transformers leveraging 32 input patches of size 64$\times$64. Still, the number of sampling coordinates for training remains the same as RDN+LTE, which is 2304 ($48^{2}$).
6.3 Real-world SR
For the real-world SR task, we introduce ESC-Real with a basic configuration of 64, 10, and 5, which denote $C$, $N$, and $M$, respectively. ESC-Real employs the same upsampler as RealESRGAN [47] and SwinIR-Real [28], utilizing it as $U$, and incorporates three layers of convolutions (c128k1g1c128k7g128c64k1g1) as $S$. Here, c, k and g denote channel, kernel, and group size for convolution, respectively. In this approach, $F^{s}$ is added into $F$. We use the RealESRGAN degradation model and DF2KOST [46] dataset to generate low-quality images. ESC-real is first trained for 400K iterations using L1 loss, then trained for 400K iterations with L1 loss, adversarial loss, and perceptual loss, using weight factors of 1, 0.1, and 1, respectively. The network architectures used for calculating adversarial loss and perceptual loss are the same as those used in RealESRGAN. In both phases, 48 patches of size 64$\times$64 are used as input for training.
7 Additional Results on Arbitrary-scale SR
This section exhibits additional results on the arbitrary-scale SR task. Additional quantitative results are measures on the four commonly used evaluation datasets, including Set5 [2], Set14 [52], B100 [33], and Urban100 [20]. Following previous research [25], we measure Peak Signal-to-Noise Ratio (PSNR) on the Y channel after cropping images’ boundary equivalent to the upscaling factor and converting them to YCbCr color space. Our ESC+LTE outperforms other methods on Urban100 at both seen and unseen scales, as shown in Table 6.
Table 6: Quantitative comparison on arbitrary-scale SR task employing LTE [25] as upsampler. The best result on PSNR is bolded.
Methods Set5 Set14 B100 Urban100
Seen Unseen Seen Unseen Seen Unseen Seen Unseen
$\times$2 $\times$3 $\times$4 $\times$6 $\times$8 $\times$2 $\times$3 $\times$4 $\times$6 $\times$8 $\times$2 $\times$3 $\times$4 $\times$6 $\times$8 $\times$2 $\times$3 $\times$4 $\times$6 $\times$8
RDN+LTE [58] 38.23 34.72 32.61 29.32 27.26 34.09 30.58 28.88 26.71 25.16 32.36 29.30 27.77 26.01 24.95 33.04 28.97 26.81 24.28 22.88
ATD-lt+LTE§ [54] 38.28 34.73 32.57 29.21 27.22 34.14 30.64 28.91 26.67 25.21 32.35 29.30 27.77 26.01 24.95 33.12 29.06 26.95 24.41 23.00
HiT-SRF+LTE§ [56] 38.27 34.74 32.59 29.25 27.21 34.03 30.64 28.91 26.68 25.21 32.37 29.30 27.76 26.00 24.94 33.18 29.05 26.89 24.34 22.94
ESC+LTE (Ours) 38.29 34.79 32.72 29.18 27.24 34.05 30.69 28.94 26.70 25.24 32.38 29.32 27.77 26.00 24.96 33.30 29.21 27.04 24.44 23.03